# X-ray panorama of the SS 433/W50 complex by SRG/eROSITA

### v1

@@ 5.2.4 Extended X-ray jets
In contrast to this “thermal” picture, the X-ray spectra of X-ray emission from the “extended X-ray jets” demonstrate a featureless spectrum that can be described by a power law with the slope in the range from (at their base, e.g., Region 5 in Fig. 5) to (at the termination points, e.g. Region 2 in Fig. 5). In agreement with previous spectral studies of this emission (e.g., Brinkmann et al., 2007), the thermal bremsstrahlung model also provides an equally good fit with temperatures from keV (at the termination) to keV (at the base). The gas emissivity required to produce the X-ray luminosity at level implies hydrogen number densities cm-3 in these regions (assuming a filling factor of unity). Clearly, if the thermal nature of this emission were true, it would have a drastically higher pressure compared to the ambient soft X-ray emitting gas and would expand sideways supersonically, driving a strong shock wave. Alternatively, the emission might result from the population of relativistic particles accelerated within EXJs. The detection of polarization in X-rays (Kaaret et al., 2024) and TeV emission from regions cospatial with the EXJs (Abeysekara et al., 2018; Alfaro et al., 2024; H. E. S. S. Collaboration et al., 2024) strongly supports the non-thermal scenario.
The absence of any noticeable radio, infrared, or optical emission from these structures would then require substantial self-absorption of radiation at these wavelengths (which is unlikely given that the gas number density is likely very low inside them) or a rather peculiar energy distribution of the relativistic particles with a large fraction of them having energies in excess of tens of TeV. Another possibility could be scattering of very bright and beamed emission of the central source (e.g., Panferov & Kabanov, 2013; Khabibullin & Sazonov, 2016, 2019), but the small optical depth with respect to the Thomson scattering, , and the observed spectral gradient along these structures disfavor this possibility.
@@ 6 Discussion
The Eastern EXJ was well mapped in X-rays with XMM-Newton, Chandra, and NuSTAR (e.g., Brinkmann et al., 2007; Safi-Harb et al., 2022), revealing its non-thermal nature. Furthermore, recent IXPE observation (Kaaret et al., 2024) revealed polarized X-ray emission from a patch adjacent to the base of the Eastern EXJ. These data suggest that X-ray photons are produced by a synchrotron mechanism with the magnetic field predominantly aligned with the EXJ axis. The eROSITA data show that the bases of both EXJs are located at essentially the same distances from SS 433, have similarly sharp inner boundaries, and non-thermal (featureless) spectra across their entire extents.
@@
These structures have often been considered as a direct consequence of the relativistic jets’ propagation inside the nebula before they deposit the bulk of their kinetic energy in the lobes, energizing them and the whole extent of the nebula (e.g., Panferov, 2014). However, the direction of the SS 433 jets is known to precess about the mean axis with the amplitude degrees, which is a factor of 2 larger than the observed opening angle of EXJ. Since this extended X-ray emission comes from distances of tens of pc, i.e. orders of magnitude further away from the central source than the compact radio and X-ray jets, which are traced up to a few cm, the actual connection between them is not obvious (e.g., Panferov, 2017). It might be governed by recollimation of the jets’ flows at some intermediate distance from the source (Eichler, 1983; Peter & Eichler, 1993), a historical change in the jets’ direction (Goodall et al., 2011a; Monceau-Baroux et al., 2015), or a different regime of baryonic jets propagation through the magnetized ISM after their gas “freezes out” after recombining due to rapid adiabatic expansion, cooling losses and lack of ambient interactions (Churazov et al., 2020).
@@
If we assume that there were no recent drastic changes in the properties of W50/SS 433 (e.g., temporary quenching of the central source), the overall picture of the “large-scale energy flow” in the W50/SS 433, supported by X-ray data, can be schematically decomposed into three major components (illustrated in Fig. 8):
@@
“Dark flow”: from a compact source up to the base of EXJs covering the range of distances from to from SS 433. Over this distance range, there are no visible signatures in the X-ray (or any other) band.
@@
“Non-thermal flow”: from the base to the end of EXJs that have featureless non-thermal spectra (presumably, synchrotron emission).
@@
“Thermal flow”: extended diffuse regions between the W50 radio boundaries and EXJs. Its X-ray emission is typical for ISM shock-heated to sub-keV temperatures.
@@
Analytical and numerical models primarily aimed at reproducing W50’s radio morphology often assume that a quasi-spherical part of the nebula is a remnant of the supernova explosion, while the elongated structures are directly associated with the energy release by the binary system (e.g., Begelman et al., 1980; Zealey et al., 1980; Eichler, 1983; Peter & Eichler, 1993; Velázquez & Raga, 2000; Zavala et al., 2008; Goodall et al., 2011a; Asahina et al., 2014; Monceau-Baroux et al., 2015; Panferov, 2017; Ohmura et al., 2021). In the model considered by Churazov et al. (2024), the entire nebula is powered by the central source. There, an anisotropic outflow/wind from the Hyper-Eddington accretor is assumed - a combination of a more powerful collimated outflow along the orbital axis and a quasi-isotropic wind in all other directions. In this model, the “Dark flow” phase corresponds to the free expansion of the wind, and the onset of the EXJs is associated with the termination shock of the isotropic wind, which initiates shocks in the collimated flow where the pressure rises suddenly. Finally, the thermal part is the shock-heated ISM. In this model, the characteristic shape of the W50 nebula (a quasi-spherical part and two extensions) reflects the anisotropy of the velocity and kinetic power of the wind, with no need for the impact of the narrow baryonic jets at all.
@@
In all these models, the outer shock, which delineates the radio boundary of W50, is not dissimilar to the shocks seen around middle-aged supernova remnants or stellar wind-blown bubbles (e.g., Konigl, 1983; Dubner et al., 1998). What makes W50 “special” is the presence of EXJs, which are radio-faint but bright in X-rays and VHE gamma-rays. These structures are also plausibly related to efficient particle acceleration at strong shocks, but with a shock configuration that is rather different compared to typical SNR shocks (e.g. Bykov et al., 2025). Remarkable symmetry of the EXJs, very clearly demonstrated by the eROSITA image, points towards irrelevance of the ambient ISM properties to the energy and momentum propagation, as well as the particle acceleration regime, meaning that a cocoon or a wind-blown cavity likely encompasses these structures in the current epoch. In this regard, the presence of H emitting filaments in several places across the nebula might help probe the energetic content of the nebula (e.g., Boumis et al., 2007; Abolmasov et al., 2010; Farnes et al., 2017). Interestingly, the location of the brightest H filaments does not correspond to any prominent features in thermal X-ray emission. This might be the case when the H emission comes from the outer boundary of the nebula, where the shock wave is already in the radiative mode. Significantly higher gas temperature obtained for the emission from the Western lobe might also be indicative of this process, namely that the radiative shock wave runs against the Galactic density gradient and stalls earlier in this direction, evacuating a smaller volume to be filled by the hot X-ray gas. Given the presence of bright silicon lines in the spectra extracted from these regions, an observation of them with the XRISM observatory (Tashiro et al., 2025) might provide estimates of the directed and turbulent velocities. A large grasp soft X-ray microcalorimeter mission, like LEM (Kraft et al., 2022) or HUBS (Bregman et al., 2023) would also be perfectly suited for in-depth spectral exploration of this complex system with multiple interacting components.
@@ 7 Conclusions
SRG/eROSITA observations of W50/SS 433 provide, for the first time, the entire X-ray map of the nebula with high spatial and spectral resolution. The data support a physical picture in which the anisotropic energy flow from SS 433 evolves passing through three distinct stages:
an invisible “dark” flow between 0.1 and 25 pc (=anisotropic wind from the binary system)
a “non-thermal” flow over another (=EXJs)
a thermal flow (= shock-heated ISM) that envelopes EXJs.
@@
The thermal part of the W50 X-ray emission can be reasonably well described by the shock-heated plasma that has not yet reached temperature and ionization equilibrium. Such emission is typical for middle-age or old SNRs. The outer radio boundary of the nebula is also reminiscent of the SNR shocks.
On the contrary, the “Extended X-ray Jets” are the most remarkable features of this system on scales of tens of pc. Their sharp inner edges plausibly correspond to extreme shocks that accelerate particles and power the X-ray (synchrotron) emission, and, at 10 orders of magnitude higher energies, the TeV emission. W50/SS 433 system illustrates clearly the important role the Hyper-Eddington accretors might play for the energetics of the ISM in galaxies at different redshifts and the production of ultra-high-energy particles.
@@ Acknowledgements.
This work is partly based on observations with the eROSITA telescope onboard SRG space observatory. The SRG observatory was built by Roskosmos in the interests of the Russian Academy of Sciences, represented by its Space Research Institute (IKI) in the framework of the Russian Federal Space Program, with the participation of the Deutsches Zentrum für Luft- und Raumfahrt (DLR). The eROSITA X-ray telescope was built by a consortium of German Institutes led by MPE, and supported by DLR. The SRG spacecraft was designed, built, launched, and operated by the Lavochkin Association and its subcontractors. The science data are downlinked via the Deep Space Network Antennae in Bear Lakes, Ussurijsk, and Baikonur, funded by Roskosmos. The development and construction of the eROSITA X-ray instrument was led by MPE, with contributions from the Dr. Karl Remeis Observatory Bamberg ECAP (FAU Erlangen-Nuernberg), the University of Hamburg Observatory, the Leibniz Institute for Astrophysics Potsdam (AIP), and the Institute for Astronomy and Astrophysics of the University of Tübingen, with the support of DLR and the Max Planck Society. The Argelander Institute for Astronomy of the University of Bonn and the Ludwig Maximilians Universität München also participated in the science preparation for eROSITA. The eROSITA data were processed using the eSASS/NRTA software system developed by the German eROSITA consortium and analyzed using proprietary data reduction software developed by the Russian eROSITA Consortium. IK acknowledges support by the COMPLEX project from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program grant agreement ERC-2019-AdG 882679.References
- Abell & Margon (1979) Abell, G. O. & Margon, B. 1979, Nature, 279, 701
- Abeysekara et al. (2018) Abeysekara, A. U., Albert, A., Alfaro, R., et al. 2018, Nature, 562, 82
- Abolmasov et al. (2010) Abolmasov, P., Maryeva, O., & Burenkov, A. N. 2010, AN, 331, 412
- Alfaro et al. (2024) Alfaro, R., Alvarez, C., Arteaga-Velázquez, J. C., et al. 2024, ApJ, 976, 30
- Arnason et al. (2021) Arnason, R. M., Papei, H., Barmby, P., Bahramian, A., & Gorski, M. D. 2021, MNRAS, 502, 5455

### v2

@@ 5.2.4 Extended X-ray jets
In contrast to this “thermal” picture, the X-ray spectra of X-ray emission from the EXJs demonstrate a featureless spectrum that can be described by a power law with a slope ranging from (at their base; e.g., Region 5 in Fig. 5) to (at the termination points; e.g. Region 2 in Fig. 5). This picture is consistent with the results of focused spectral studies of the western EXJ, including the data above 10 keV from NuSTAR observations (Safi-Harb et al., 2022). Also, in agreement with previous spectral studies of this emission (e.g., Brinkmann et al., 2007), the thermal bremsstrahlung model also provides an equally good fit with temperatures from keV (at the termination) to keV (at the base). The gas emissivity required to produce the X-ray luminosity at level implies hydrogen number densities of cm-3 in these regions (assuming a filling factor of unity). Clearly, if the thermal nature of this emission were true, it would have a drastically higher pressure compared to the ambient, soft X-ray-emitting gas and would expand sideways supersonically, driving a strong shock wave. Alternatively, the emission might result from the population of relativistic particles accelerated within EXJs. The detection of polarization in X-rays (Kaaret et al., 2024) and TeV emission from regions cospatial with the EXJs (Abeysekara et al., 2018; Alfaro et al., 2024; H. E. S. S. Collaboration et al., 2024) strongly supports the nonthermal scenario.
The absence of any noticeable radio, infrared, or optical emission from these structures would then require substantial self-absorption of radiation at these wavelengths (which is unlikely given that the gas number density is likely very low inside them) or a rather peculiar energy distribution of the relativistic particles with a large fraction of them having energies in excess of tens of TeV. Another possibility could be scattering of very bright and beamed emission of the central source (e.g., Panferov and Kabanov, 2013; Khabibullin and Sazonov, 2016, 2019), but the small optical depth with respect to the Thomson scattering, , and the observed spectral gradient along these structures disfavor this possibility.
@@ 6 Discussion
The eastern EXJ was well mapped in X-rays with XMM-Newton, Chandra, and NuSTAR (e.g., Brinkmann et al., 2007; Safi-Harb et al., 2022), revealing its nonthermal nature. Furthermore, recent IXPE observation (Kaaret et al., 2024) revealed polarized X-ray emission from a patch adjacent to the base of the eastern EXJ. These data suggest that X-ray photons are produced by a synchrotron mechanism with the magnetic field predominantly aligned with the EXJ axis. The eROSITA data show that the bases of both EXJs are located at essentially the same distances from SS 433 and have similarly sharp inner boundaries and nonthermal (featureless) spectra across their entire extents.
@@
These structures have often been considered as a direct consequence of the relativistic jets’ propagation inside the nebula before they deposit the bulk of their kinetic energy in the lobes, energizing them and the whole extent of the nebula (e.g., Panferov, 2014). However, the direction of the SS 433 jets is known to precess at the mean axis with an amplitude of degrees, which is a factor of two larger than the observed opening angle of an EXJ. Since this extended X-ray emission comes from distances of approximately tens of parsec; i.e., orders of magnitude further away from the central source than the compact radio and X-ray jets, which are traced up to a few cm, the actual connection between them is not obvious (e.g., Panferov, 2017). It might be governed by recollimation of the jets’ flows at a given intermediate distance from the source (Eichler, 1983; Peter and Eichler, 1993), a historical change in the jets’ direction (Goodall et al., 2011a; Monceau-Baroux et al., 2015), or a different regime of baryonic jet propagation through the magnetized interstellar medium (ISM) after their gas “freezes out” following recombination due to rapid adiabatic expansion, cooling losses, and lack of ambient interactions (Churazov et al., 2020).
@@
If we assume that there were no recent drastic changes in the properties of W50/SS 433 (e.g., temporary quenching of the central source), the overall picture of the “large-scale energy flow” in W50/SS 433, supported by X-ray data, can be schematically decomposed into the three major components (illustrated in Fig. 8) listed below.
@@
Dark flow: this spreads from a compact source up to the base of EXJs covering the range of distances from to from SS 433. Over this distance range, there are no visible signatures in the X-ray (or any other) band.
@@
Nonthermal flow: this is from the base to the end of EXJs that have featureless nonthermal spectra (presumably synchrotron emission).
@@
Thermal flow: this represents extended diffuse regions between the W50 radio boundaries and EXJs. Its X-ray emission is typical for an ISM that is shock-heated to sub-keV temperatures.
@@
Analytical and numerical models primarily aimed at reproducing W50’s radio morphology often assume that a quasi-spherical part of the nebula is a remnant of the supernova explosion, while the elongated structures are directly associated with the energy release by the binary system (e.g., Begelman et al., 1980; Zealey et al., 1980; Eichler, 1983; Peter and Eichler, 1993; Velázquez and Raga, 2000; Zavala et al., 2008; Goodall et al., 2011a; Asahina et al., 2014; Monceau-Baroux et al., 2015; Panferov, 2017; Ohmura et al., 2021). In the model considered by Churazov et al. (2024), the entire nebula is powered by the central source. There, an anisotropic outflow/wind from the hyper-Eddington accretor is assumed: a combination of a more powerful collimated outflow along the orbital axis and a quasi-isotropic wind in all other directions. In this model, the darkflow phase corresponds to the free expansion of the wind, and the onset of the EXJs is associated with the termination shock of the isotropic wind, which initiates shocks in the collimated flow where the pressure rises suddenly. Finally, the thermal part is the shock-heated ISM. In this model, the characteristic shape of the W50 nebula (a quasi-spherical part and two extensions) reflects the anisotropy of the velocity and kinetic power of the wind, with no need for the impact of the narrow baryonic jets at all.
@@
In all these models, the outer shock, which delineates the radio boundary of W50, is not dissimilar to the shocks seen around middle-aged SNRs or stellar wind-blown bubbles (e.g., Konigl, 1983; Dubner et al., 1998). What makes W50 “special” is the presence of EXJs, which are radio-faint but bright in X-rays and VHE gamma-rays. These structures are also plausibly related to efficient particle acceleration at strong shocks, but with a shock configuration that is rather different compared to typical SNR shocks (e.g. Bykov et al., 2025) . The remarkable symmetry of the EXJs – which is very clearly demonstrated by the eROSITA image– points toward the irrelevance of the ambient ISM properties to the energy and momentum propagation, as well as the particle acceleration regime, meaning that a cocoon or a wind-blown cavity likely encompasses these structures in the current epoch. In this regard, the presence of H-emitting filaments in several places across the nebula might help probe the energetic content of the nebula (e.g., Boumis et al., 2007; Abolmasov et al., 2010; Farnes et al., 2017). Interestingly, the location of the brightest H filaments does not correspond to any prominent features in thermal X-ray emission. This might be the case when the H emission comes from the outer boundary of the nebula, where the shock wave is already in the radiative mode. Significantly higher gas temperature obtained for the emission from the western lobe might also be indicative of this process, namely that the radiative shock wave runs against the Galactic density gradient and stalls earlier in this direction, evacuating a smaller volume to be filled by the hot X-ray gas. Given the presence of bright silicon lines in the spectra extracted from these regions, an observation of them with the XRISM observatory (Tashiro et al., 2025) might provide estimates of the directed and turbulent velocities. A large grasp soft X-ray microcalorimeter mission such as LEM (Kraft et al., 2022) or HUBS (Bregman et al., 2023) would also be perfectly suited for the in-depth spectral exploration of this complex system with multiple interacting components.
@@ 7 Conclusions
The SRG/eROSITA observations of W50/SS 433 provide, for the first time, the entire X-ray map of the nebula with high spatial and spectral resolution. The data support a physical picture in which the anisotropic energy flow from SS 433 evolves passing through the following three distinct stages:
an invisible dark flow between 0.1 and 25 pc (=anisotropic wind from the binary system)
a nonthermal flow over another (=EXJs)
a thermal flow (= shock-heated ISM) that envelopes EXJs
@@
The thermal part of the W50 X-ray emission can be reasonably well described by the shock-heated plasma that has not yet reached temperature and ionization equilibrium. Such emission is typical for middle-aged or old SNRs. The outer radio boundary of the nebula is also reminiscent of the SNR shocks.
On the contrary, the EXJs are the most remarkable features of this system on scales of tens of parsec . Their sharp inner edges plausibly correspond to extreme shocks that accelerate particles and power the X-ray (synchrotron) emission, and, at energy levels that are ten orders of magnitude higher, the TeV emission. The W50/SS 433 system clearly illustrates the important role the hyper-Eddington accretors might play in the energetics of the ISM in galaxies at different redshifts and the production of ultra-high-energy particles.
Data availability
X-ray data analyzed in this article were used with the permission of the Russian SRG/eROSITA consortium. The reduced broad-band image (Fig. 2) is available at the CDS via anonymous ftp to cdsarc.cds.unistra.fr (130.79.128.5).
@@ Acknowledgements.
This work is partly based on observations with the eROSITA telescope onboard SRG space observatory. The SRG observatory was built by Roskosmos in the interests of the Russian Academy of Sciences, represented by its Space Research Institute (IKI) in the framework of the Russian Federal Space Program, with the participation of the Deutsches Zentrum für Luft- und Raumfahrt (DLR). The eROSITA X-ray telescope was built by a consortium of German Institutes led by MPE, and supported by DLR. The SRG spacecraft was designed, built, launched, and operated by the Lavochkin Association and its subcontractors. The science data are downlinked via the Deep Space Network Antennae in Bear Lakes, Ussurijsk, and Baikonur, funded by Roskosmos. The development and construction of the eROSITA X-ray instrument was led by MPE, with contributions from the Dr. Karl Remeis Observatory Bamberg ECAP (FAU Erlangen-Nuernberg), the University of Hamburg Observatory, the Leibniz Institute for Astrophysics Potsdam (AIP), and the Institute for Astronomy and Astrophysics of the University of Tübingen, with the support of DLR and the Max Planck Society. The Argelander Institute for Astronomy of the University of Bonn and the Ludwig Maximilians Universität München also participated in the science preparation for eROSITA. The eROSITA data were processed using the eSASS/NRTA software system developed by the German eROSITA consortium and analyzed using proprietary data reduction software developed by the Russian eROSITA Consortium. IK acknowledges support by the COMPLEX project from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program grant agreement ERC-2019-AdG 882679. Some of the maps were built using cubehelix colorscheme developed by David Green (Green, 2011).References
- A kinematic model for SS433. Nature 279 (5715), pp. 701–703. External Links: Document, ADS entry Cited by: §1.
- Very-high-energy particle acceleration powered by the jets of the microquasar SS 433. Nature 562 (7725), pp. 82–85. External Links: Document, 1810.01892, ADS entry Cited by: §1, §5.2.4.
- The Eastern filament of W50. AN 331 (4), pp. 412. External Links: Document, 1003.4706, ADS entry Cited by: §1, §6.
- Spectral Study of Very-high-energy Gamma Rays from SS 433 with HAWC. ApJ 976 (1), pp. 30. External Links: Document, 2410.21796, ADS entry Cited by: §1, §5.2.4.
- Distances to Galactic X-ray binaries with Gaia DR2. MNRAS 502 (4), pp. 5455–5470. External Links: Document, 2102.02615, ADS entry Cited by: footnote 1.